OpenStudy (anonymous):

How would I figure out what equation of a line that contains the origin and the point (1,2).?

OpenStudy (anonymous):

y=mx+b where m is the slope and b is the y intercept Origin: (0,0) Point: (1,2) Slope: \[(y _{2}-y _{1})/(x _{2}-x _{1})=m\] 2-0/1-0=m 2/1=m 2=m b=0 y=2x

OpenStudy (anonymous):

y= mx+b is suppose to be a formula.?

OpenStudy (primeralph):